a woman can do a particular job in 10 hour and her daughter can do it in 12 hour. After a woman and her daughter have been working for 1 hour, they are joined by the older daughter. The three completed the job in 3 hour. How long will it take the older daughter to do the job herself?
Answer by mananth(16949)

woman---- 10------0.10
daughter------- 12-------0.08
Together they do 0.18 of the job in one hour

After 1 hour
Balance job 0.82
let second daughter take x hours to do 1 job
so she does 1/x of the job in 1 hour.
All three do (1/x + 0.18 ) of the job in 1 hour
they have to complete 0.82 of the job so they will take
0.82/(1/x+0.18) hours ==> 3
0.82/(1/x+0.18)= 3
0.82= 3(1/x+0.18)
0.82/3 -0.18 = 1/x
1/x= 0.093
x= 10.75 hours.